OpenStudy (anonymous):

Medals given Describe two ways to estimate the product 5 and one-eighth times 7 and nine-tenths. Find the estimate for this product using each method.

OpenStudy (austinl):

1/8 = 0.125 9/10 = 0.9 7+5=12 To estimate, you could say ~13 This is one way.

OpenStudy (austinl):

Not sure how to estimate another way, perhaps wait for another helper?

OpenStudy (anonymous):

ok